OpenStudy (yusmin):

Use guess and check to find how many years it will take for your money to double if you have 2,000 invested with a growth rate of 7% per year, rounded up to the nearest year.

OpenStudy (yusmin):

The money will double after how many years?

OpenStudy (dumbcow):

If we are just estimating then there is a quick rule to use. Divide rate into 72 to get num of years it takes to double your initial capital. ---> 72/7 = 10.3 So about 10 years

OpenStudy (yusmin):

ohhhh hmm yeah I have no idea what I was doing I got something totally off lol

OpenStudy (dumbcow):

For more exact answer, solve following equation \[1.07^n = 2\] \[n = \frac{\ln 2}{\ln 1.07}\]

OpenStudy (yusmin):

thank you so much

OpenStudy (photon336):

\[[A_{0}]*(1.07)^{n} = 2[A_{0}]\] and then solve for n